This word problem is really bugging me can you solve it?
A father is now 20 years older than his son. In 8 years the father will be 5 years more than twice the sons age. Find their present age

Answer by mananth(12270)

A father is now 20 years older than his son. In 8 years the father will be 5 years more than twice the sons age. Find their present age
...
son = x
father = x+20
...
after 8 years
son = x+8
father x+20+8=x+28
..
x+28 = 2(x+8)+5
x+28=2x+16+5
x+28=2x+21
2x-x=28-21
x= 7 son's age now.
Father's age now = x+20 = 27 years
...
m.ananth@hotmail.ca